Calculate the number of grams of aluminum sulfate that could sulfate be obtained by the action of 12.5 grams of aluminum on a excess of sulfuric acid.


2 Al + 3H2SO4 --> Al2(SO4)3 + 3H2


1
Expert's answer
2021-09-09T09:47:03-0400

Molar masses are:

Al: 26.98 g/mol;

Al2(SO4)3: 342.15 g/mol.


Therefore,

"m(Al_2(SO_4)_3)=12.5g(Al)\times\frac{1mol(Al)}{26.98g(Al)}\times\frac{1mol(Al_2(SO_4)_3)}{2mol(Al)}\times\frac{342.15g(Al_2(SO_4)_3)}{1mol(Al_2(SO_4)_3)}=79.3g"


Answer: 79.3 g